Copper has a resistivity of Ω m. A wire of diameter 1.5 mm and length 25 m is connected across a potential difference of 50 V.

Calculate:

(a) The resistance of the wire

(b) The current

(c) The power dissipated in the wire
1
Expert's answer
2020-11-30T14:53:57-0500

Resistivity of copper is "1.7\times10^{-8}"


Wire diameter"=1.5mm \implies radius=7.5\times10^{-4}m"

Wire length"=25m"


Voltage applied across"=50V"


a) Resistance of the wire"(R) =\rho\frac{l}{a}=1.7\times10^{-8}\frac{25}{\pi \times(7.5\times10^{-4}) ^2}"


"R=0.2405\varOmega"


b) Current"(I) =\frac{V}{R}=\frac{50}{0.2405}=207.9A"

c) Power dissipated"=I^2R=(207.9) ^2\times 0.2405=10394.989W"